OpenStudy (anonymous):

How do i expand these expressions? please help.? log2(a^6b^4) & Ln(√x/y^3)

OpenStudy (anonymous):

write without exponents.

OpenStudy (saifoo.khan):

\[\large \frac{loga^2b^4}{\log2}\]

OpenStudy (saifoo.khan):

we have to write it without expoenents that's the main question...

OpenStudy (anonymous):

yes the question is . Expand the following expressions. Write without exponents.

OpenStudy (saifoo.khan):

i know a rule, it's: \[\Large \log_a b^c = c \log_ab\]

OpenStudy (saifoo.khan):

\[\frac{6 \log(a) + 4 \log(b)}{\log 2}\]

OpenStudy (anonymous):

so thats the answer? thanks. Appreciate it.

OpenStudy (saifoo.khan):

Yes welcime.

OpenStudy (saifoo.khan):

welcome*

OpenStudy (anonymous):

do you know how to do the second one?

OpenStudy (saifoo.khan):

\[\Large \frac12 (\log (x) - 6 \log(y))\]